OpenStudy (anonymous):

The acceleration function (m/s^2) and the initial velocity are given for a particle moving along a line. Find each of the following: a(t)=2t+3, v(0)=-4, 0 is less than or equal to t which is less than or equal to 3 A)the velocity at time t B) The distance traveled during the given time interval. C) Show that for motion in a straight line with constant acceleration a, and initial velocity v0, and initial displacement s0, the displacement

OpenStudy (freckles):

s'=v v'=a so to find v you integrate a

OpenStudy (freckles):

then to find the constant of integration there use v(0)=-4

OpenStudy (amistre64):

is this a calculus based question?

OpenStudy (amistre64):

 the velocity is the area of the acceleration graph, if this is not calculus based

OpenStudy (anonymous):

I was able to figure out parts a and b, I mostly need help with c

OpenStudy (anonymous):

it's calculus based

OpenStudy (amistre64):

your C part is missing things ...

OpenStudy (amistre64):

let a(t) = a integrate to find velocity integrate again to fine position

OpenStudy (anonymous):

oh whoops the full c says "Show that for motion in a straight line with constant acceleration a, initial velocity v0, and initial displacement s0, the displacement after time t is S=1/2at^2+v0t+s0

OpenStudy (amistre64):

what do we get when we integrate a constant? \[\int a~d=a\int dt = ??\]

OpenStudy (amistre64):

... dropped a t at the start lol

OpenStudy (anonymous):

you would get t right?

OpenStudy (anonymous):

no wait i mean you would get ad

OpenStudy (amistre64):

a (t+c) for some arbitraty constant c v = at + ac, but ac is just still an arbitrary constant v = at + c, what is c when t=0?

OpenStudy (amistre64):

\[v =\int a~dt=at+c\] \[v_0=a(0)+c\] therefore \[v = at+v_0\] integrate it again ....

OpenStudy (amistre64):

\[s=\int at+v_o~dt\] \[s=\int at~dt+\int v_o~dt\] \[s=a\int t~dt+v_0\int ~dt\]

OpenStudy (amistre64):

thers no real ned to split it, just work it how you feel best